Robert Limited has just paid a current (annual) dividend of $1 which is expected to grow by 20% at the end of one year from now and then by 15% at the end of two years from now. The dividend growth rate will stabilize at a constant 5% p.a. thereafter. If the cost of equity for Robert Limited is 20% p.a., what is the fair value of a Robert Limited share at the present time?
Answer #1
Year Dividend PVF @20% PV of dividend
1 1.2 0.833 1
2 1.38 0.694 0.9583
PV 1.9583

Dividend at the end of year 2 = 1.38* 105% = 1.449

Price after year 2 = Dividend / (Cost of equity - growth rate )

= 1.449 / (20% -5%)

=9.66

Present Value of year 2 Stock = 9.66 * 0.694 = 6.708

Fair value of the share = 6.708 + 1.9583 =$ 8.67